OpenStudy (anonymous):

help help please . please Which statement is true concerning the graph of two lines showing the given rates? Rate 1: 12 mi in 2 h Rate 2: 15 mi in 3 h A. The graph with Rate 1 passes through (2, 6). B. The graph with Rate 2 passes through (3, 5). C. The graph with Rate 1 passes through (6, 1). D. The graph with Rate 2 passes through (2, 10).

OpenStudy (anonymous):

For rate one you can write a function f(h) = 12/2 = 6h and for the second one g(h) = 15/3 = 5h The only answer that checks out is D. since g(2) = 5 * 2= 10

OpenStudy (anonymous):

thanx